OpenStudy (anonymous):

solve the initial value problem: dy/dx - y/x = xe^x , y(1) = e - 1 I know the answer is y = xe^x - x i just dont know the steps to get to it

OpenStudy (jamesj):

The integrating factor is 1/x So (1/x)y' + (-1/x^2)y = e^x i.e., [ (1/x)y ]' = e^x thus (1/x)y = e^x + c => y = x.e^x + cx Now apply your initial condition.